OpenStudy (anonymous):

The sum of a number and 12 times its reciprocal is 8. Find the number.

OpenStudy (anonymous):

\[x + \frac{ 12 }{ x } = 8\]

OpenStudy (anonymous):

multiply everything by x and solve

OpenStudy (anonymous):

so 3/2

OpenStudy (anonymous):

or do i use the quadratic formula

OpenStudy (anonymous):

you need quadratic formula or any other method (like completing square)

OpenStudy (anonymous):

for x^2 -8x + 12 = 0

OpenStudy (anonymous):

you can also factorize it to (x-2)(x-6) = 0

OpenStudy (anonymous):

so is it 6,2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks